OpenStudy (anonymous):

Would someone expand this using mclaurins? (x+2)/(x−4)(x−2)

OpenStudy (anonymous):

at x = 0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use partial fractions first

OpenStudy (anonymous):

you need to split it up first

OpenStudy (anonymous):

yeah i did that. 2/(2-x) - 3/(4-x)

OpenStudy (anonymous):

A long division will give the McLauren series for the function.

OpenStudy (anonymous):

then factor out the constants and use geometric series sum

OpenStudy (anonymous):

better idea than what i was going to do for sure.

OpenStudy (anonymous):

\[= \frac{ 2}{ 2 ( 1 - \frac{x}{2} ) } - \frac{3}{ 4 ( 1- \frac{x}{4} ) } \]

OpenStudy (anonymous):

stucked after partial fraction. :( I dont want to do mclaurins separately.

OpenStudy (anonymous):

\[=\sum_{n=0}^{\infty}[ (\frac{x}{2} ) ^n - \frac{3}{4} (\frac{x}{4})^n ]\]

OpenStudy (anonymous):

you can simplify the sum a bit if you want

OpenStudy (anonymous):

shouldn't the constant be 1/4? if i put n = 0 i get -1/4 here? not to interrupt because this all looks good

OpenStudy (anonymous):

\[=\sum_{n=0}^{\infty} (1- \frac{ 3}{2^{n+2}} ) (\frac{x}{2})^n \]

OpenStudy (anonymous):

think thats it, something similair